A start-up company needs to raise $5 million for a project. It plans to issue 22-year bonds with a par value of $1,000 and a coupon rate of 6.7% with interest paid seminannually. How many approximate bonds will the company need to issue to raise the $5 million if investors demand a required return of 5%? Enter your answer as a whole number (to the nearest integer).
Answer #1

Information provided:

Par value= future value= $1,000

Time= 22 years*2= 44 semi-annual periods

Coupon rate= 6.7%/2= 3.35%

Coupon payment= 0.0335*1,000= $33.50

Required return= 5%/2= 2.50% per semi-annual period

The question is solved by first calculating the price of the bonds.

The present value is calculated to compute the price of the bonds.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

PMT= 33.50

N= 44

I/Y= 2.50

Press the CPT key and PV to compute the present value.

The value obtained is 1,225.28.

Therefore, the price of the bonds is $1,225.28.

The number of bonds needed to be issued= $5,000,000/ $1,225.28

                                                                                 = 4,080.70 bonds\rightarrow    4,081 bonds.

In case of any query, kindly comment on the solution.